### v1

@@ Abstract
To solve ever more complex problems, Deep Neural Networks are scaled to billions of parameters, leading to huge computational costs. An effective approach to reduce computational requirements and increase efficiency is to prune unnecessary components of these often over-parameterized networks. Previous work has shown that attribution methods from the field of eXplainable AI serve as effective means to extract and prune the least relevant network components in a few-shot fashion. We extend the current state by proposing to explicitly optimize hyperparameters of attribution methods for the task of pruning, and further include transformer-based networks in our analysis. Our approach yields higher model compression rates of large transformer- and convolutional architectures (VGG, ResNet, ViT) compared to previous works, while still attaining high performance on ImageNet classification tasks. Here, our experiments indicate that transformers have a higher degree of over-parameterization compared to convolutional neural networks. Code is available at https://github.com/erfanhatefi/Pruning-by-eXplaining-in-PyTorch.
Keywords:
Explainable AI Pruning Attribution Optimization
@@ 1 Introduction
In recent years, Deep Neural Networks (DNNs) have been growing increasingly larger, demanding more computational resources and memory. To address these challenges, several efficient architectures, such as MobileNet [23] or EfficientFormer [29], have been proposed to reduce computational costs. However, the gain in efficiency comes at the cost of performance, as inherently efficient architectures struggle to keep pace with the recent surge in high-performing transformer models.
The key challenge in pruning (without sequential re-training) is the identification of structures within the model that can be removed without adversely affecting its performance. Earlier works have shown that structures may be chosen according to, e.g., parameter magnitudes [19, 28], functional redundancy [17], sensitivity [48, 32], or importance in the decision-making process [49, 7]. Especially the latter approach has been promising, utilizing tools from the seemingly unrelated field of XAI that originally seeks to explain the model reasoning process to human stakeholders.
@@
While prior attribution-based pruning [49] achieved remarkable results, it is restricted to heuristically chosen LRP hyperparameters and limited to specific Convolutional Neural Networks (CNNs) and toy models.
Following up upon the work [35] that optimized LRP w.r.t. specific explainability metrics, we argue that LRP or any other attribution method can be optimized specifically for the goal of more effective pruning (see Fig. 1). In addition, recent works [2, 4] have extended LRP to transformer-based architectures, allowing us now to also apply attribution-based pruning to the Vision Transformer (ViT) architecture.
@@
    revealing that attributions optimized for explanations are not necessarily the best for pruning. Notably, simple hyperparameter settings show high performance across all tested architectures.
@@ 2 Related Work
Explainability and Local Feature Attribution Methods
Research in local explainability led to a plethora of methods [30, 37], commonly resulting in local feature attributions quantifying the importance of input features in the decision-making process. These attributions are often shown in the form of heatmaps in the vision domain. Notably, methods based on (modified) gradients, backpropagate attributions from the output to the input through the network, conveniently offering attributions of all latent components and neurons in a single backward pass [44, 42]. Gradient-based attribution methods, however, can suffer from noisy gradients, rendering them unreliable for deep architectures [6]. Prominently, LRP [5, 33] introduces a set of different rules (with hyperparameters) that allow to reduce the noise level. In fact, as shown in [35], attribution methods such as LRP can be optimized for certain XAI criteria, e.g., faithfulness or complexity [33, 22]. We follow up on this observation, and specifically optimize XAI w.r.t. pruning.
Pruning of Deep Neural Networks
For pruning CNNs, either individual (kernel) weights or whole structures, e.g., filters of convolution layers or neurons can be recognized as candidates for pruning [21]. For transformer architectures, such structures include heads of attention modules or linear layers inside transformer blocks [46, 27]. In order to prune such structures, several criteria have been proposed to indicate which components are best suited to be removed, retaining performance as best as possible. The work of [19] suggests pruning parameters based on weight magnitudes, offering a computationally-free criterion. Alternatively, in [11], the authors propose to prune neurons based on their activation patterns. However, recent work [7] has shown that weight or activation magnitudes do not necessarily reflect a component’s contribution during the inference process, e.g., also a small weight has the potential to be very relevant for the predication of a class.
Explainability For Efficient Deep Learning
The work of [49] introduces a novel pruning criterion based on XAI, and proposes to use latent relevance values extracted from the attribution method of LRP [5, 34] to assign importance scores to network structures. By taking into account how structures are used during the inference process, [49] can effectively improve the efficiency-performance tradeoff. Later, the works of [7, 43] also illustrate the value of XAI methods for network quantization. In all these works, heuristically chosen hyperparameters for LRP have been used, which overlooks the potential for specific optimization to the task of pruning. Notably, LRP is model-specific and thus restricted to compatible architectures. In this work, we also include recent LRP extensions to transformer architectures, and observe, that simple LRP settings commonly result in high pruning performances.
@@ 3.1 Attribution-based Pruning
For our structured pruning framework, we view a DNN as a collection of $p$ (interlinked) components $\Psi=\{\psi_{1},\dots,\psi_{p}\}$, that can correspond to, e.g., whole layers, (groups of) neurons, convolutional filters or attention heads. We further assume access to an attribution method that generates attribution scores (relevance) $R_{\psi_{k}}({x_{i}})$ of component $\psi_{k}\in\Psi$ for the prediction of a sample $x_{i}$. The overall relevance of $\psi_{k}$ is then estimated through the mean relevance over a set of reference samples $\mathcal{X}_{\text{ref}}=\{x_{1},x_{2},\dots,x_{n_{\text{{ref}}}}\}$ as
Figure 2: Attribution-based pruning workflow: Firstly, the relevant model structures are found by explaining a set of reference samples. The attribution method of choice (here LRP) highlights the components and paths in the network which positively and negatively contribute to the decision-making. Positive and negative relevances are indicated by red and blue color respectively, and components with white color indicate no relevance. Removing structures that receive the least relevance results in a sparser subnetwork, which performs significantly better than after random pruning. Notably, relevances can be computed w.r.t. a subset of output classes (e.g., “corgi” only), resulting in a subnetwork specifically designed to perform the restricted task. Credit: Nataba/iStock.
$\bar{R}_{\psi_{k}}=\frac{1}{n_{\text{ref}}}\sum_{i=1}^{n_{\text{ref}}}R_{\psi_{k}}({x_{i}})\,.$ (1)
Ultimately, we collect relevances for all components via the set $\mathcal{R}$, given as
=$\mathcal{R}=\{\bar{R}_{\psi_{1}},\bar{R}_{\psi_{2}},\dots,\bar{R}_{\psi_{p}}\}\,,$ (2)
@@
where 1 is an indicator function with condition $i\in\{c\}_{q}$. The whole attribution-based pruning workflow is depicted in Fig. 2. Notably, previous work [49] demonstrates that LRP is an effective method for attributing latent structures, as used for Eq. 1. LRP further offers several hyperparameters to tune, which makes it a versatile choice for our framework.
@@ 3.2 Layer-wise Relevance Propagation
Beginning with an initial relevance score $R_{j}^{L}$ at output $j$ of layer $f^{L}$, the score is layer-by-layer redistributed to its input variables depending on the contribution from the input variables to the output value.
@@
$\centering R_{i\leftarrow j}^{(l-1,l)}=\frac{z_{ij}}{z_{j}}R_{j}^{l}\,.\@add@centering$ (6)
@@
$R_{i}^{l-1}=\sum_{j}R_{i\leftarrow j}^{(l-1,l)}.$ (7)
This process ensures the property of relevance conservation between adjacent layers:
@@
When a group of neurons performs the same task within a convolutional channel or attention head, it is beneficial to aggregate the total relevance of the entire group into a single relevance score. This aggregation process helps in simplifying the analysis and interpretation of the model’s behavior by focusing on the collective relevance of each convolutional channel or attention head, rather than examining individual neurons. Further discussions of component-wise aggregation are given in Appendix 0.B.
@@ 3.3 Tuneable Hyperparameters of LRP
Within the LRP framework, various rules (as detailed in Sec. 0.A.2) have been proposed to tune relevance computation for higher explainability.
Further, several works have shown that using different rules for different architecture parts (referred to as a “composite”) [33, 25, 2] is favorable for explanations as it improves robustness and faithfulness. We follow up on this idea, and perform a large-scale analysis over a variety of rules for different network parts.
Inspired by the tuning settings of [35], we split a network into four parts:
@@
For each group, a specific rule can be chosen, which together ultimately form the LRP composite. We refrain from treating each individual layer as a separate group due to the increased computational expenses associated with this setting.
For transformer architectures, the works of [2, 4] proposed novel LRP rules to attribute softmax non-linearities in the attention modules (see Sec. 0.A.2 for more details), encouraging us to also optimize the LRP rule for the softmax non-linearities, resulting in an additional hyperparameter here.
It is important to note, that relevance values can be positive or negative. Components with positive relevance values indicate that they significantly contribute towards the class prediction. Conversely, components with negative relevance values diminish the classification score, effectively speaking against the predicted class, as visible in Fig. 2 where “cat” features (and the respective model components) speak against the “corgi” prediction. Components with near zero relevance do not contribute to the inference process in any meaningful way. Thus, a question arises w.r.t. the optimal order of pruning: Should we start with negative, or near zero relevant components? To that end, we also add a hyperparameter indicating whether absolute relevance values are used as in Eq. 1, i.e., $\bar{R}_{\psi_{k}}\mapsto|\bar{R}_{\psi_{k}}|$ for all $\psi_{k}\in\Psi$ if true. The work of [49] takes into account only positive attributions by their LRP rule choice (see LRP-$z^{+}$ in Sec. 0.A.2). Our experiments in Appendix 0.F, however, reveal that the highest pruning rates are achieved by first pruning components with near zero relevance, as their minimal contribution ensures low impact on the overall model performance.
@@ 3.4 Hyperparameter Optimization Procedure
where the performance as given by $C$ depends on the network $f$ and its pruned parameters $\Psi_{\text{PR}_{i}}(\theta)$ after pruning step $i$ using the hyperparameter setting $\theta$ for the attribution method. Ultimately, our objective will be the maximization of $A_{\text{PR}}$ w.r.t. $\theta$. Hyperparameter search is performed via grid search and Bayesian optimization techniques, as detailed in Appendix 0.D.
@@ 4.1 How Over-Parameterized are Vision Models?
Figure 3: Investigating over-parameterization in DNNs through attribution-based pruning (LRP-$\epsilon$ in blue, unoptimized in orange color) and random pruning (green color). We compare pruning of all models w.r.t. different task difficulties, i.e., to differentiate between 1000 (dotted line), 100 (dashed line) or three ImageNet classes (solid line). High performance for high sparsification rates indicates over-parameterization, i.e., many network components are not important for the task. Compared to the ResNet-18 and VGG-16-BN CNNs (top), the ViT-B-16 transformer shows a higher degree of over-parameterization (bottom). SEM is illustrated (shaded area) in the current and all other figures.
@@
For baseline comparisons, we utilize random pruning, Integrated Gradient [44] and our optimized LRP attributions (see Sec. 4.2). We intentionally exclude naive weight magnitude or activation pruning since [49] has previously demonstrated that XAI-based methods outperform these traditional approaches. Notably, especially for simpler tasks, attribution-based pruning outperforms random pruning, as attributions are output-specific and allow identifying the sub-graph that is relevant for the restricted task, as also illustrated in Fig. 2.
Compared to the previously discussed CNNs, the ViT-B-16 vision transformer seems to be more over-parameterized, as visible in Fig. 3 (bottom). Up to a pruning rate of 20 %, no meaningful accuracy loss is measured for detecting the 1000 ImageNet classes, irrespective of whether attention heads or neurons in linear layers are pruned. This might be due to the fact, that the vision transformer consists of about 46 thousand neurons, which is approximately ten times the number of convolutional filters of the CNNs. Also, when comparing the ResNet-50 and ResNet-18 (see Fig. A.1 for ResNet-50), the larger network shows higher pruning potential.
@@ 4.2 Finding the Optimal Attributions for CNNs
Figure 4: Pruning CNNs models pre-trained on ImageNet (simplified task to detecting three classes), using ten reference samples per class. Results show a better sparsification-performance tradeoff for our optimized LRP composite compared to a heuristic (faithful) LRP composite, Yeom et al. [49] (details for each in Tab. A.2, Sec. 0.A.5, and Sec. 0.A.3.3) and random pruning.
@@
Whereas we begin with optimizing LRP for CNN pruning, we follow up with transformers in the next section. Regarding CNNs, we can strongly reduce the accuracy-sparsity tradeoff when pruning the convolution filters with our method based on optimizing LRP attributions compared to other baselines, e.g., the heuristically chosen variant of LRP used by [49] (see Sec. 0.A.3.3), as also illustrated in Fig. 4 and Tab. 1. Extensive hyperparameter search reveals that across different CNN architectures, simple hyperparameter settings exist that are well-suited for pruning in general. One such candidate is the LRP-$\epsilon$ rule (see Eq. A.3) applied for all convolutional layers. We refer to Tab. A.2 for a list of the best performing hyperparameter settings. Interestingly for CNNs, LRP composites known to lead to faithful explanations (in the suggested context of [38, 5]) optimized by [25], are also effective in pruning. One such composite (detailed in Sec. 0.A.5) is also shown in Fig. 4 and denoted as “Faithful LRP”. Moreover, our results indicate a larger amount of unused structures across architectures with shortcut connections, i.e., the ResNet models. As intermediate features can be passed through the short-cuts, layers have the potential to be not used at all.

### v2

@@ Abstract
To solve ever more complex problems, Deep Neural Networks are scaled to billions of parameters, leading to huge computational costs. An effective approach to reduce computational requirements and increase efficiency is to prune unnecessary components of these often over-parameterized networks. Previous work has shown that attribution methods from the field of eXplainable AI serve as effective means to extract and prune the least relevant network components in a few-shot fashion. We extend the current state by proposing to explicitly optimize hyperparameters of attribution methods for the task of pruning, and further include transformer-based networks in our analysis. Our approach yields higher model compression rates of large transformer and convolutional architectures (VGG, ResNet, ViT) compared to previous works, while still attaining high performance on ImageNet classification tasks. Here, our experiments indicate that transformers have a higher degree of over-parameterization compared to convolutional neural networks. Code is available at https://github.com/erfanhatefi/Pruning-by-eXplaining-in-PyTorch.
Keywords Explainable AI  $\cdot$ Pruning  $\cdot$ Attribution Optimization
@@ 1 Introduction
In recent years, Deep Neural Networks (DNNs) have been growing larger increasingly, demanding ever more computational resources and memory. To address these challenges, several efficient architectures, such as MobileNet [22] or EfficientFormer [28], have been proposed to reduce computational costs. However, the gain in efficiency comes at the cost of performance, as inherently efficient architectures struggle to keep pace with the recent surge in high-performing but costly transformer models.
The key challenge in pruning (without sequential re-training) is the identification of structures within the model that can be removed without adversely affecting its performance. Earlier works have shown that structures may be chosen according to, e.g., parameter magnitudes [18, 27], functional redundancy [16], sensitivity [47, 31], or importance in the decision-making process [48, 6]. Especially the latter approach has been promising, utilizing tools from the seemingly unrelated field of eXplainable Artificial Intelligence (XAI) that originally seeks to explain the model reasoning process to human stakeholders.
@@
While prior attribution-based pruning [48] achieved remarkable results, it is restricted to heuristically chosen LRP hyperparameters and specific Convolutional Neural Networks (CNNs).
Following up upon the work [34] that optimized LRP w.r.t. specific explainability metrics, we argue that LRP or any other tunable attribution method can be optimized specifically for the goal of more effective pruning (see Fig. 1). In addition, recent works [2, 3] have extended LRP to transformer-based architectures, allowing us now to also apply attribution-based pruning to the Vision Transformer (ViT) architecture.
@@
    revealing that attributions optimized for explanations are not necessarily the best for pruning.
@@ 2 Related Work
Explainability and Local Feature Attribution Methods.
Research in local explainability led to a plethora of methods (e.g., [29, 36]), commonly resulting in local feature attributions quantifying the importance of input features in the decision-making process. These attributions are often shown in the form of heatmaps in the vision domain. Notably, methods based on (modified) gradients, backpropagate attributions from the output to the input through the network, conveniently offering attributions of all latent components and neurons in a single backward pass [43, 41]. Gradient-based attribution methods, however, can suffer from noisy gradients, rendering them unreliable for deep architectures [5]. Prominently, LRP [4, 32] introduces a set of different rules (with hyperparameters) that allow to reduce the noise level. In fact, as shown in [34], attribution methods such as LRP can be optimized for certain XAI criteria, e.g., faithfulness or complexity [32, 21]. We follow up on this observation, and specifically optimize XAI w.r.t. the task of pruning. That is, we add Neural Network (NN) pruning as an XAI evaluation criterion to optimize for.
Pruning of Deep Neural Networks.
For pruning CNNs, either individual (kernel) weights or whole structures, e.g., filters of convolution layers or neurons can be recognized as candidates for pruning [20]. For transformer architectures, such structures include heads of attention modules or linear layers inside transformer blocks [45, 26]. In order to prune such structures, several criteria have been proposed to indicate which components are best suited to be removed, retaining performance as best as possible. The work of [18] suggests pruning parameters based on weight magnitudes, offering a computationally free criterion. Alternatively, in [10], the authors propose to prune neurons based on their activation patterns. However, recent work [6] has shown that weight or activation magnitudes do not necessarily reflect a component’s contribution during the inference process, e.g., also a small weight has the potential to be very relevant for the prediction of a class.
Explainability For Efficient Deep Learning.
The work of [48] introduces a novel pruning criterion based on XAI, and proposes to use latent relevance values extracted from the attribution method of LRP [4, 33] to assign importance scores to network structures. By taking into account how structures are used during the inference process, the work of [48] can effectively improve the efficiency-performance tradeoff. Later, the works of [6, 42] also illustrate the value of XAI methods for network quantization. In all these works, heuristically chosen hyperparameters for LRP have been used, which overlooks the potential for specific optimization to the task of pruning. Notably, LRP is model-specific and thus restricted to compatible architectures. In this work, we also include recent LRP extensions to transformer architectures, and observe, that a specific LRP rule commonly result in high pruning performances.
@@ 3.1 Attribution-based Pruning
For our structured pruning framework, we view a DNN as a collection of $p$ (interlinked) components $\Psi=\{\psi_{1},\dots,\psi_{p}\}$, that can correspond to, e.g., whole layers, (groups of) neurons, convolutional filters or attention heads. We further assume access to an attribution method that generates attribution scores (relevance scores) $R_{\psi_{k}}({x_{i}})$ of component $\psi_{k}\in\Psi$ for the prediction of a sample $x_{i}$. The overall relevance of $\psi_{k}$ is then estimated through the mean relevance over a set of reference samples $\mathcal{X}_{\text{ref}}=\{x_{1},x_{2},\dots,x_{n_{\text{{ref}}}}\}$ as
Figure 2: Attribution-based pruning workflow: Firstly, the relevant model structures are identified by explaining a set of reference samples. The attribution method of choice (here LRP) highlights the components and paths in the network which positively and negatively contribute to the decision-making. Positive and negative relevances are indicated by red and blue color respectively, and components with white color indicate zero or low relevance. Removing structures that receive the least relevance results in a sparser subnetwork, which performs significantly better than after random pruning. Notably, relevances can be computed w.r.t. a subset of output classes (e.g., “corgi” only), resulting in a subnetwork specifically designed to perform the restricted task. Credit: Nataba/iStock.
$\bar{R}_{\psi_{k}}=\frac{1}{n_{\text{ref}}}\sum_{i=1}^{n_{\text{ref}}}R_{\psi_{k}}({x_{i}})$ (1)
We collect relevance scores for all components via the set $\mathcal{R}$, given as
=$\mathcal{R}=\{\bar{R}_{\psi_{1}},\bar{R}_{\psi_{2}},\dots,\bar{R}_{\psi_{p}}\}$ (2)
@@
where 1 is an indicator function with condition $i\in\{c\}_{q}$. The whole attribution-based pruning workflow is depicted in Fig. 2. Notably, previous work [48] demonstrates that LRP is an effective method for attributing latent structures. LRP further offers several hyperparameters to tune, which makes it a versatile choice for our framework.
@@ 3.2 Layer-wise Relevance Propagation
Beginning with an initial relevance score $R_{j}^{L}$ at output $j$ of layer $f^{L}$ (usually set as $f_{j}^{L}$ for an output of choice $j$), the score is layer-by-layer redistributed through all latent structures to its input variables depending on the contribution from the these units to the output value:
@@
$\centering R_{i\leftarrow j}^{(l-1,l)}=\frac{z_{ij}}{z_{j}}R_{j}^{l}\@add@centering$ (6)
@@
$R_{i}^{l-1}=\sum_{j}R_{i\leftarrow j}^{(l-1,l)}$ (7)
This process ensures relevance conservation between adjacent layers:
@@
When a group of neurons performs the same task as within a convolutional channel or attention head, it is beneficial to aggregate the total relevance of the entire group into a single relevance score. This aggregation process helps in simplifying the analysis and interpretation of the model’s behavior by focusing on the collective relevance of each convolutional channel or attention head, rather than examining individual neurons. Further discussions of component-wise aggregation are given in Appendix B.
@@ 3.3 Tuneable Hyperparameters of LRP
Within the LRP framework, various rules (as detailed in Sec. A.2) have been proposed to tune relevance computation for higher explainability.
Further, several works have shown that applying different rules to variable parts of a network (referred to as a “composite”) [32, 24, 2] enhances faithfulness and robustness of explanations. We follow up on this idea, and perform a large-scale analysis over a variety of rules for different network parts. Inspired by the tuning configurations of [34], we split a network into four parts:
@@
For each group, a specific rule set can be chosen, which together ultimately form the LRP composite. We refrain from treating each individual layer as a separate group due to the increased computational expenses associated with this setting during hyperparameter search.
For transformer architectures, the works of [2, 3] proposed novel LRP rules to attribute softmax non-linearities in the attention modules (see Sec. A.2 for more details), encouraging us to also optimize LRP configurations for the softmax non-linearities, resulting in an additional hyperparameter choices here.
It is important to note, that relevance values can be positive or negative. Components with positive relevance values indicate that they significantly contribute towards the explained class prediction. Conversely, components with negative relevance values diminish the classification score, effectively speaking against the predicted class, as visible in Fig. 2 where “cat” features (and the respective model components) speak against the “corgi” prediction. Components with near zero relevance do not contribute to the inference process in any meaningful way. Thus, a question arises w.r.t. the optimal order of pruning: Should we start with negative, or near zero relevant components? To that end, we also add a hyperparameter indicating whether absolute relevance values are used as in Eq. 1, i.e., $\bar{R}_{\psi_{k}}\mapsto|\bar{R}_{\psi_{k}}|$ for all $\psi_{k}\in\Psi$ if true. The work of [48] takes into account and computes only positive attributions by their LRP rule choice (see LRP-$z^{+}$ in Sec. A.2). Our experiments in Appendix F, however, reveal that the highest pruning rates are achieved by computing both positive and negative relevance signals, and then first pruning components with near zero relevance, as their minimal contribution in any direction ensures low impact on the overall model performance.
@@ 3.4 Hyperparameter Optimization Procedure
where the performance as given by $C$ depends on the network $f$ and its pruned parameters $\Psi_{\text{PR}_{i}}(\theta)$ after pruning step $i$ using the hyperparameter setting $\theta$ for the attribution method. Ultimately, our objective will be the maximization of $A_{\text{PR}}$ w.r.t. $\theta$. Hyperparameter search is performed via grid search and Bayesian optimization techniques, as detailed in Appendix D.
@@ 4.1 How Over-Parameterized are Vision Models?
Figure 3: Investigating over-parameterization in DNNs through attribution-based pruning (LRP-opt in blue, Integrated Gradient in orange color) and random pruning (green color). We compare pruning of all models w.r.t. different task difficulties, i.e., to differentiate between 1000 (dotted line), 100 (dashed line) or three ImageNet classes (solid line). High performance for high sparsification rates indicates over-parameterization, i.e., many network components are not important for the task. Compared to the ResNet-18 and VGG-16-BN CNNs (top), the ViT-B-16 transformer shows a higher degree of over-parameterization (bottom). SEM is illustrated (shaded area) in the current and all other figures.
@@
For baseline comparisons, we utilize random pruning, Integrated Gradient [43] and our optimized LRP attribution (LRP-opt) (see Sec. 4.2). We intentionally exclude naive weight magnitude or activation pruning since [48] has previously demonstrated that XAI-based methods outperform these traditional approaches by a large margin. Notably, especially for simpler tasks, attribution-based pruning outperforms random pruning, as attributions are output-specific and allow identifying the sub-graph that is relevant for the restricted task, as also illustrated in Fig. 2.
Compared to the previously discussed CNNs, the ViT-B-16 vision transformer seems to be more over-parameterized, as visible in Fig. 3 (bottom), and therefore easier to prune. Up to a pruning rate of 20 %, no meaningful accuracy loss is measured for detecting the 1000 ImageNet classes, irrespective of whether attention heads or neurons in linear layers are pruned. This might be due to the fact, that the vision transformer consists of about 46,000 neurons, which is approximately ten times the number of convolutional filters of the CNNs. Also, when comparing the ResNet-50 and ResNet-18 (see Fig. A.1 for ResNet-50), the larger network shows higher pruning potential.
@@ 4.2 Finding the Optimal Attributions for CNNs
Figure 4: Pruning CNNs models pre-trained on ImageNet (simplified task to detecting three classes), using ten reference samples per class. Results show a better sparsification-performance tradeoff for our optimized LRP composite compared to a heuristic (faithful) LRP composite, Yeom et al. [48] (details for each in Tab. A.2, Sec. A.5, and Sec. A.3.3) and random pruning.
@@
Whereas we begin with optimizing LRP for CNN pruning, we follow up with transformers in the next section. Regarding CNNs, we can strongly reduce the accuracy-sparsity tradeoff when pruning the convolution filters with our method based on optimizing LRP attributions compared to other baselines, e.g., the heuristically chosen variant of LRP used by [48] (see Sec. A.3.3), as also illustrated in Fig. 4 and Tab. 1. Extensive hyperparameter search reveals that across different CNN architectures, simple hyperparameter settings exist that are well-suited for pruning in general. One such candidate is the LRP-$\epsilon$ rule (see Eq. A.3) applied for all convolutional layers. We refer to Tab. A.2 for a list of the best performing hyperparameter settings. Interestingly for CNNs, LRP composites known to lead to faithful explanations (in the suggested context of [37, 4]) optimized by [24], are also effective in pruning. One such composite (detailed in Sec. A.5) is also shown in Fig. 4 and denoted as “Faithful LRP”. Moreover, our results indicate a larger amount of unused structures across architectures with shortcut connections, i.e., the ResNet models. As intermediate features can be passed through the short-cuts, while layers have the potential to be not used at all.